# Incremental Sequence Labeling: A Tale of Two Shifts

### v1

@@ Appendix B Datasets
Table 3: Examples of inputs and outputs for each dataset.
@@
Table 4: Detailed description of each dataset.
@@
FL means the full ground-truth label for all steps. During the learning process, we will label unseen entities as non-entities [O].
Appendix C Baselines
@@ Appendix E Additional Experimental Results
The previous method exhibits a rapid decrease in probability distribution with increasing incremental steps, coinciding with a decline in the F1 score. In contrast, our approach IS3 effectively mitigates the model’s penalization of old entities, thereby maintaining good performance.
@@
# Time (Min) # Trainable Params each Task
SEQ 150 109M
SelfTrain 276 109M
ExtendNER 182 109M
CFNER 512 109M
DLD 158 109M
RDP 188 109M
OCILNER 420 109M
ICE 126 28K
CPFD 282 109M
Ours 155 109M
Table 7: Comparisons with state-of-the-art methods on OntoNotes5. The best results are highlighted in bold and the second best results are underlined.
@@
((a)) i2b2 (FG-1-PG-1)
((b)) i2b2 (FG-2-PG-2)
((c)) i2b2 (FG-8-PG-1)
((d)) i2b2 (FG-8-PG-2)
((e)) OntoNotes5 (FG-1-PG-1)
((f)) OntoNotes5 (FG-2-PG-2)
((g)) OntoNotes5 (FG-8-PG-1)
((h)) OntoNotes5 (FG-8-PG-2)
((a)) ExtendNER
((b)) IS3 (Ours)

### v2

@@ 6 Conclusion
Acknowledgements
The work described in this paper was partially funded by the National Natural Science Foundation of China (Grant No. 62272173), the Natural Science Foundation of Guangdong Province (Grant Nos. 2024A1515010089, 2022A1515010179), the Science and Technology Planning Project of Guangdong Province (Grant No. 2023A0505050106), and the National Key R&D Program of China (Grant No. 2023YFA1011601).
@@ Appendix B Datasets
Table 3 shows the detailed description of each dataset. Table 4 shows the format of inputs and outputs for Sequence labeling task. FL means the full ground-truth label for all steps. During the learning process, we will label unseen entities as non-entities [O].
Appendix C Baselines
    Table 3: Detailed description of each dataset.
@@ Appendix C Baselines
    Table 4: Examples of inputs and outputs for each dataset.
@@ Appendix E Additional Experimental Results
As shown in Figure 11, the previous method exhibits a rapid decrease in probability distribution with increasing incremental steps, coinciding with a decline in the F1 score. In contrast, our approach IS3 effectively mitigates the model’s penalization of old entities, thereby maintaining good performance.
# Time (Min) # Trainable Params each Task
SEQ 150 109M
SelfTrain 276 109M
ExtendNER 182 109M
CFNER 512 109M
DLD 158 109M
RDP 188 109M
OCILNER 420 109M
ICE 126 28K
CPFD 282 109M
IS3(Ours) 155 109M
@@
We compare the runtime and the number of updated parameters in Table 5. The results indicate that the proposed IS3 requires less than 5% more training time than SEQ and less training time than most previous methods, such as CFNER and CPFD.
Table 7: Comparisons with state-of-the-art methods on the OntoNotes5 dataset using the bert-base-cased model. The best results are highlighted in bold and the second best results are underlined.
@@
Table 8: Comparisons with state-of-the-art methods on the i2b2 dataset using the roberta-base model. The best results are highlighted in bold and the second best results are underlined. The average of each incremental step is provided in Figure 5.
Dataset Methods FG-1-PG-1 FG-2-PG-2 FG-8-PG-1 FG-8-PG-2
i2b2 FT 2.68 ± 1.19 14.36 ± 0.81 7.39 ± 1.58 23.02 ± 0.42 1.91 ± 0.38 15.56 ± 1.77 6.00 ± 0.49 25.01 ± 0.73
SelfTrain 17.58 ± 1.60 37.95 ± 1.10 25.84 ± 3.26 43.80 ± 3.24 6.97 ± 1.15 28.73 ± 0.48 27.85 ± 3.72 45.80 ± 2.37
ExtendNER 17.68 ± 1.75 34.53 ± 2.58 26.84 ± 1.63 44.33 ± 3.21 9.49 ± 1.07 28.80 ± 0.72 22.27 ± 7.08 40.07 ± 5.62
CFNER 32.65 ± 1.87 47.06 ± 2.70 43.12 ± 2.84 54.61 ± 2.40 33.52 ± 0.78 38.61 ± 1.56 36.19 ± 7.72 49.46 ± 6.09
DLD 16.26 ± 4.79 34.02 ± 3.46 27.12 ± 7.30 45.81 ± 3.26 5.54 ± 1.68 27.80 ± 1.39 20.39 ± 4.46 38.43 ± 2.58
RDP 21.70 ± 0.88 40.71 ± 2.86 35.38 ± 2.79 53.81 ± 0.98 26.49 ± 6.16 39.54 ± 3.96 42.98 ± 4.06 56.25 ± 1.44
OCILNER 9.27 ± 4.60 27.98 ± 1.94 15.14 ± 9.85 34.44 ± 4.11 15.89 ± 8.39 35.54 ± 7.28 21.26 ± 8.70 39.21 ± 5.01
ICE_PLO 29.89 ± 0.23 36.65 ± 0.32 32.14 ± 0.36 37.34 ± 0.31 34.34 ± 0.84 41.07 ± 0.91 39.88 ± 0.71 44.34 ± 0.46
ICE_O 25.56 ± 0.94 35.57 ± 0.75 33.17 ± 0.27 38.16 ± 0.25 32.77 ± 1.16 40.58 ± 1.30 36.95 ± 1.41 41.95 ± 1.44
CPFD 13.65 ± 4.16 40.81 ± 3.10 26.38 ± 1.72 49.82 ± 1.41 4.59 ± 0.86 31.76 ± 2.39 25.23 ± 9.73 48.85 ± 3.37
IS3 (Ours) 34.14 ± 1.45 51.61 ± 1.60 47.91 ± 2.76 59.22 ± 1.01 42.99 ± 3.01 52.25 ± 1.96 52.36 ± 3.42 60.99 ± 1.49
Table 9: Comparisons with state-of-the-art methods on the OntoNotes5 dataset using the roberta-base model. The best results are highlighted in bold and the second best results are underlined.
Dataset Methods FG-1-PG-1 FG-2-PG-2 FG-8-PG-1 FG-8-PG-2
OntoNotes5 FT 1.75 ± 0.11 12.43 ± 0.11 5.07 ± 0.32 20.85 ± 0.27 1.32 ± 0.26 12.03 ± 0.20 4.97 ± 0.33 20.82 ± 0.14
SelfTrain 38.88 ± 6.38 47.57 ± 2.30 54.02 ± 0.76 55.78 ± 1.87 39.82 ± 5.23 49.05 ± 2.34 51.77 ± 2.28 57.71 ± 0.48
ExtendNER 23.29 ± 5.15 36.77 ± 6.26 44.00 ± 2.82 50.06 ± 1.44 24.12 ± 1.85 40.39 ± 3.00 44.48 ± 5.62 53.50 ± 2.86
CFNER 41.86 ± 2.78 48.73 ± 3.11 54.24 ± 1.04 58.07 ± 1.71 51.51 ± 2.06 55.65 ± 0.56 52.23 ± 2.75 58.95 ± 4.11
DLD 25.25 ± 4.17 37.24 ± 6.06 46.58 ± 1.78 50.93 ± 0.54 24.04 ± 2.54 40.63 ± 0.91 45.89 ± 1.27 54.89 ± 0.38
RDP 34.84 ± 2.92 45.90 ± 0.91 46.95 ± 1.49 55.02 ± 0.73 41.58 ± 3.37 53.69 ± 0.53 52.54 ± 0.44 62.03 ± 1.05
OCILNER 27.98 ± 5.18 29.86 ± 4.28 29.64 ± 4.27 37.24 ± 2.79 29.94 ± 1.60 42.51 ± 1.13 30.90 ± 2.41 45.54 ± 0.63
ICE_PLO 33.66 ± 0.66 37.21 ± 0.36 36.17 ± 0.07 37.71 ± 0.59 36.90 ± 0.19 40.33 ± 0.37 38.83 ± 0.34 41.10 ± 0.30
ICE_O 30.46 ± 1.69 33.41 ± 1.91 35.05 ± 0.92 37.17 ± 1.12 37.38 ± 1.14 40.22 ± 0.83 40.69 ± 0.21 42.25 ± 0.47
CPFD 29.50 ± 1.47 43.12 ± 2.18 41.01 ± 2.92 51.56 ± 0.52 40.58 ± 2.50 51.64 ± 2.52 50.70 ± 3.65 61.14 ± 1.57
IS3 (Ours) 42.69 ± 1.21 49.73 ± 1.13 58.77 ± 1.97 59.27 ± 1.28 55.43 ± 2.01 60.25 ± 1.49 61.75 ± 0.96 64.61 ± 0.78